Hi,I am struggling with this question,please explain:
The bond order in O2+ Is the same as:
C2/O2 2-/F2-/N2 2+ Or B2
Thanks
1
Expert's answer
2020-05-01T14:31:04-0400

Bond order is defined as (nb - na)/2, where nb = number of electrons on bonding molecular orbitals, na = number of electrons on antibonding molecular orbitals.

O2+ has 8 electrons on bonding orbitals (two electrons on each of sigma-s, sigma-p, and two pi-p orbitals) and 3 electrons on the antibonding simga*-s and pi*-p orbitals, so the bond order is 2.5.

For the same reasoning and counting the electrons, one obtains that the bond orders (BO) in the other particles are:

C2 BO = 2

O22- BO = 1

F2- BO = 0.5

N22+ BO = 2

B2  BO = 1

None of them fits the bond order of O2+